How would you simplify 5/6+2/3*3/4-1/4*2/5 to equal 37/30? I tried it and got 82/30.
This is how I did it:
5/6+2/3*3/4-1/4*2/5
I canceled the threes out and got 1*2 then I also cross cancelled the 2 and 4 making it a 1 and 2, I did that again to the 1/4 and the 2/5 leaving me with this.
5/6+2-1/10
after that I found the lcd which is 30
I Plug that in leaving me with this
5/6=25/30
2/1=60/30
1/10=3/30
when you plug those in to the equation they equal 82/30
the answer is suppose to be 37/30

Answer by swincher4391(1107)

Let's add some parenthesis to emphasize order of operations here.
5/6+(2/3*3/4)-(1/4*2/5)
Left most parenthesis:
(2/3) * (3/4) cancel out the threes to get (2/4) which reduces to 1/2.
Right most parenthesis
(1/4) * (2/5) cancel out the 2, reduce the 4 to a 2 to get (1/(2*5)) = 1/10.
so we have (5/6) + (1/2) - (1/10)
Least Common Multiple of 6,12 and 10 is 60.
(50/60) + (30/60) - (6/60) = 74/60 (reduces to) 37/30 <---- Answer